# A high resolution simulation of protoplanetary disk turbulence driven by the vertical shear instability

### v1

@@ 4.2 Midplane vorticity
Figure 12(c) shows the perturbation radial vorticity: it has the lowest values among the three and is occasionally associated with other vorticity components. It most likely arises from the tilting of $\omega_{z}$ by $\omega_{\phi}$ KH vortices.
@@ 4.3 Vorticity in a horizontal plane at two scale heights
Perturbation vorticity components increase in magnitude away from the midplane and are therefore able to withstand the stabilizing effects of rotation and shear.
@@ 4.4 Radial wavelength in 3D
Figure 13(a) shows $\delta\omega_{z}$ layers (labeled (A)–(D)) that can be used to infer a wavelength. These layers are coherent over large azimuthal extent. We find that the ratio of the interlayer spacing to the local instability wavelength is 4.4 for layers (A) and (B), 4.5 for (B) and (C), and 2.4 for (C) and (D). Figure 14 shows $\delta\omega_{z}$ centered at $\phi=90^{\circ}$: the spacing ratio for layers (E) and (F) is 1.5. Therefore the spacing ratio increases with radius. The inexorable radial wavelength growth present in axisymmetric VSI is halted in 3D and a time invariant local radial length scale is established.
@@ 4.5 Specific kinetic energy spectra in the midplane
where the hat denotes a real to complex Fourier transform, a star denotes the complex conjugate, and $m$ is the azimuthal wavenumber. A Hanning window was applied in $r$ since it is not a periodic direction. For $S_{uu}(k_{r})$, an average was taken in $\phi$, and for $S_{uu}(m)$ an average was taken in $r$. Note that spatial inhomogeneity of the flow in $r$ precludes a rigorous mathematical interpretation of $S_{uu}(k_{r})$. It is presented here mostly to identify the range of length scales resolved in $r$.
Remark. Dr. D. Melon Fuksman (private communication) has suggested that since the flow has nonuniform density, it would be more appropriate to consider the quantity
@@
Figure 16(a) displays the spectrum with respect to $k_{r}$ and shows that $u_{z}$ dominates the kinetic energy for most of the wavenumber range. A bump corresponding to energy injection by the primary VSI mode is clearly observed. This is followed by a short (about 0.7 of a decade) power law range with a slope of $\approx-1.84$. This is close to the $-2$ slope that a jump in velocity across a shear-layer would produce (Townsend & Taylor, 1951). Figure 16(b) displays spectra with respect to the azimuthal wavenumber $m$. The total energy (blue) has a slope close to the Kolmogorov value of $-5/3$. However, there is a small dip (relative to a straight line) at $m\approx 40$; we see that it arises from the radial component (purple line). Given that $n_{\phi}=1024$, the location of the dip corresponds to a wavelength of about 25 grid points. Again, $u_{z}$ (red) is the dominant component though not as strongly as was the case for $k_{r}$ spectra. Radial velocity fluctuations (violet) are not much smaller than vertical velocity fluctuations (red) in the inertial (power law) range. Azimuthal velocity fluctuations (green) have the least amplitude throughout the wavenumber range.
@@ 4.7 Reynolds stresses
was found to be almost constant at $\approx 0.2$-$0.3$ throughout the domain. This suggests that the shapes of the structures responsible for $T_{r\phi}$ are similar at all locations of the domain; only their intensity and/or size varies. The trace $T_{kk}$ (not shown), which is twice the kinetic energy, peaks at the midplane but maintains high values up to $z\approx 1.5H_{0}$. Strong peaks in all the stresses near the left and right radial boundary are considered to be artifacts due to boundary conditions.
@@ 4.8 The turbulence $\alpha$ parameter
We currently lack an explanation for this increase. An explanation one might initially suggest is the fact that the appropriate velocity scale for $\nu_{\mathrm{t}}$ should be the driving vertical shear $\Delta u_{\phi}(r)=|u_{\phi}(r,z=H)-u_{\phi}(r,z=0)|$ rather than the sound speed as assumed by Shakura & Sunyaev (1973). Unfortunately, however, $\Delta u_{\phi}(r)\propto c_{\mathrm{i}}(r)$ for the VSI basic state.
Figure 19 shows that with increasing resolution (starting with the brown curve) $\alpha$ initially rapidly increases to the green curve but then decreases to the blue and black curves. Inspection of the meridional plane vorticity for the brown case reveals a longer radial wavelength than the prediction (7) of linear theory for the most amplified mode. There is also minimal radial waviness of shear layers, which would contribute to a suppression of $T_{\phi r}$. Why should the highest resolution case (black curve) have lower $\alpha(r)$ than the less resolved blue and green cases? We conjecture that large scale waviness present at lower resolution breaks down into smaller vortices at higher resolution, leading to a smaller $T_{r\phi}$ correlation.
@@ 5 Late time artifacts in the axisymmetric simulation
Given that flow length scales increase with time in our simulation (see below) as well as in the simulations of M24a, i.e., that there is an inverse cascade, the time to reach stationarity in any given simulation would depend on the domain size.
@@
Figure 23 shows the perturbation azimuthal and vertical vorticity components at $t/T_{0}=197.6$, which is at a late time but prior to the stationary state. Their structure is very different from the 3D case and what is observed at early times in the axisymmetric case. The dominant component now is $\omega_{z}$ rather than $\omega_{\phi}$, which translates to a larger magnitude of the azimuthal velocity perturbation compared to the vertical velocity. The shear layers of $\omega_{\phi}$ that formed the edges of vertical jets at earlier times have been completely disrupted, and panel (a) is reminiscent of 2D turbulence. Panel (b) shows that the $\approx 7$ layers of vertical vorticity seen at $t/T_{0}=36.1$ (Figure 5) have now intensified, and only 4 remain. At a later time ($t/T_{0}=403.7$, not shown) only 2 such layers remain. In other words, flow length scales increase with time. The layers of strong $\delta\omega_{z}$ are associated with jumps in the total specific angular momentum $j_{z}=u_{\phi}r$ (panel c), which has a strong vertically coherent staircase profile with respect to $r$.
@@
This can be seen more clearly in Figure 24, which shows the midplane profiles of $j_{z}(r)$ at $t/T_{0}=197.6$ and $t/T_{0}=403/7$. The latter is in the stationary state and has only 2 jumps in $j_{z}$. As shown earlier (Figure 4, blue line), in 3D there is no large scale rearrangement of angular momentum, rather only a weak one in the upper layers of the disk. The circulation, $\Gamma=2\pi j_{z}$, of circular material lines is proportional to $j_{z}$. Hence by Stokes’ theorem, a region of nearly constant $\Gamma(r)$ must be irrotational. For inviscid barotropic flow, Kelvin’s theorem asserts that that $D\Gamma/Dt=0$ for each circular material line. Hence there are two possible mechanisms for the formation of the staircase structure: radial mixing by eddies between the vertical jets or baroclinic torque. For a circular material line we have that
@@
where $A$ is the circular surface enclosed by the circle and $\widehat{\mathbf{z}}$ is the unit vector normal to the surface. However, the $z$-component of the baroclinic term in the integrand is zero for axisymmetric flow. Hence, we conclude that radial mixing by eddies in regions bounded by vertical jets is responsible for the staircase profile of $j_{z}$.
@@
In the axisymmetric simulations of MF24a there is also a clear difference between early and late times. Let us focus on their dg3c4_2048 simulation. Their Figure 1 (left hand plots) and Figure 2 (top) shows very many small steps in the $j_{z}$ staircase at an earlier time. On the other hand, their Figure 11 shows the same run at 1360 orbits. Clearly an increase of length scales has taken place leading to only 3 jumps in $j_{z}$.
@@ 6.1 Caveats for applying RWI to VSI simulations
following fluid particles. Here $P$ is the vertically integrated pressure, $\Sigma$ is the surface density, and $\gamma$ is the adiabatic exponent. A necessary condition for RWI is that, in the basic state, the diagnostic function
@@
has a local maximum. The thin disk assumption is valid only for motions whose horizontal length scale $\ell\gg H(r)$. On the other hand, the radial size of LPVs observed in simulations is about $H(r)$. In addition, the locally isothermal equation of state, $p=\rho c_{\mathrm{i}}^{2}(r)$, assumed in the present and other VSI simulations implies that radial motions fail to conserve $S$ following fluid particles. Likewise, it remains to be shown whether radiative hydrodynamic VSI simulations that invoke RWI (e.g., Flock et al., 2020) possess an effective adiabatic exponent. Despite the lack of adiabaticity, previous works plot a vertically averaged $\mathcal{L}(r)$, presumably setting $\gamma=1.4$. In spite of these caveats, it appears that the appearance of LPVs in VSI simulations is correlated with some (but not all) local maxima in $\mathcal{L}(r)$ and we will discuss these simulations in the next subsection.
@@ 6.4 Speculation
Why do M20 observe LPVs in the interior of the domain for the same disk parameters but we do not? The answer may lie in the link, discussed by MF24b, between maxima in $\mathcal{L}(r)$ to the $j_{z}(r)$ staircase that is prominent in axisymmetric $j_{z}(r)$. Specifically, a region of flattened $j_{z}(r)=u_{\phi}r$ corresponds to irrotational flow ($u_{\phi}\approx\mathrm{const}/r$) and therefore suppressed $\omega_{z}$. Since mass also accumulates in these regions, a local maximum is created in the first factor ($\Sigma/2\omega_{z}$) of $\mathcal{L}$. Therefore, one possibility is that M20 has a stronger $j_{z}(r)$ staircase in 3D than the present simulation, which leads to several $\mathcal{L}(r)$ maxima in the interior of the domain. It is tempting to attribute this to M20’s lower resolution; however, our low resolution simulations also failed to produce LPVs away from boundaries. Further effort is required to ascertain the reasons for this difference between our and the M20 simulation.
@@ 7 Closing remarks
    Radial wavelength. In axisymmetric VSI, the radial wavelength at early times corresponds to the most amplified linear mode, but then increases in time. In 3D, the radial wavelength is about 4.5 times larger than that of the most amplified linear mode.
@@
    Layer of positive $\delta\omega_{z}$. Axisymmetric VSI has coherent vertical layers of $\delta\omega_{z}>0$ that increase in strength with $|z|$ and are coincident with interfaces between the upward and downward jets (Figure 5). Their formation and the preference for $\delta\omega_{z}>0$ can be explained by the vertical transport of specific angular momentum $j_{z}\equiv u_{\phi}r$ by the vertical jets. This phenomenon is equivalent to the formation of a staircase profile of $j_{z}(r)$ first discovered by Klahr et al. (2023) and MF24a,b in their axisymmetric simulations. We find that these layers are also present in 3D but with much weaker jumps in specific angular momentum compared to the axisymmetric case (Figure 4(b)). These layers become stronger and more azimuthally coherent with increasing $|z|$ and undergo KH-like instability (Figure 13a), likely modified by background shear and rotation. The resulting vortices are small and possess all three vorticity components.
@@
    Variation of $\alpha$ with $r$. There is a linear increase of $\alpha(r)$ with radius with a slope (at the highest resolution) of about 14% the midradius value per scale height. The mechanism leading to this increase remains to be explained.
@@
    Late time artifacts in axisymmetric VSI. As discussed in items 1 and 2, some features of the 3D simulation can be understood by appealing to physics in the early stage of the axisymmetric simulation. However, at late times the axisymmetric simulations have artifacts that are not present in 3D. These include increasing flow length scales and a strong staircase pattern in the specific angular momentum $j_{\phi}=u_{\phi}r$ with an accompanying redistribution of mass.
We are grateful to Dr. J. David Melon Fuksman (Max-Planck-Institut für Astronomie, Heidelberg) for his detailed comments, to Drs. Paul Estrada and Ted Manning at NASA Ames for performing a careful internal review, and to the referee for many useful suggestions that led to an improved manuscript. OMU acknowledges partial support from NASA’s TCAN (Theoretical and Computational Astrophysical Networks) grant #80NSSC21K0497.

### v2

@@ 4.2 Midplane vorticity
Figure 12(c) shows the perturbation radial vorticity: it has the lowest values among the three and is occasionally associated with other vorticity components. It most likely arises from the tilting and stretching of $\omega_{z}$ by $\omega_{\phi}$ KH vortices.
@@ 4.3 Vorticity in a horizontal plane at two scale heights
Perturbation vorticity components increase in magnitude away from the midplane and are therefore able to withstand the stabilizing effects of Keplerian rotation and shear.
@@ 4.4 Radial wavelength in 3D
The $\delta\omega_{z}$ layers in Figure 13(a) (labeled (A)–(D)) can be used to infer a wavelength. These layers are coherent over a large azimuthal extent. We find that the ratio of the interlayer spacing $\ell_{r}$ to the local instability wavelength is 4.4 for layers (A) and (B), 4.5 for (B) and (C), and 2.4 for (C) and (D). Figure 14 shows $\delta\omega_{z}$ centered at $\phi=90^{\circ}$: the spacing ratio for layers (E) and (F) is 1.5. Therefore the spacing ratio increases with radius. This means that the spacing $\ell_{r}(r)$ increases with $r$ faster than $H(r)$ does. The inexorable radial wavelength growth with time present in axisymmetric VSI is halted in 3D and a time invariant local radial length scale is established. An interesting question is whether $\ell_{r}(r)$ is an appropriate length scale for determining the turbulent viscosity $\nu_{\mathrm{t}}$.
@@ 4.5 Specific kinetic energy spectra in the midplane
where the hat denotes a real to complex Fourier transform, a star denotes the complex conjugate, and $m$ is the azimuthal wavenumber. To obtain $S_{uu}(k_{r})$, a Hanning window was applied in $r$ since it is not a periodic direction, after which an average was taken in $\phi$. For $S_{uu}(m)$, an average was taken in $r$. Note that spatial inhomogeneity of the flow in $r$ precludes a rigorous mathematical interpretation of these spectra.
Remark. Dr. J. David Melon Fuksman (private communication) has suggested that since the flow has nonuniform density, it would be more appropriate to consider the quantity
@@
Figure 16(a) displays the spectrum with respect to $k_{r}$ and shows that $u_{z}$ dominates the kinetic energy for most of the wavenumber range. A bump corresponding to energy injection by VSI is clearly observed at $k_{r}H_{0}\approx 4.5$ which corresponds to a wavelength of $\ell_{r}\approx 1.4H_{0}$. On the other hand, the most linearly unstable wavelength is $\lambda_{r,\mathrm{max}}=0.31$ at the midradius of the domain. This bump is followed by a short (about 0.7 of a decade) power law range with a slope of $\approx-1.84$. This is close to the $-2$ slope that a jump in velocity across a shear-layer would produce (Townsend & Taylor, 1951). Figure 16(b) displays spectra with respect to the azimuthal wavenumber $m$. The total energy (blue) has a slope close to the Kolmogorov value of $-5/3$. However, there is a small dip (relative to a straight line) at $m\approx 40$ which corresponds to a wavelength of $\approx 9^{\circ}$; we see that the dip arises from the radial component (purple line). Again, $u_{z}$ (red) is the dominant component though not as strongly as was the case for $k_{r}$ spectra. Radial velocity fluctuations (violet) are not much smaller than vertical velocity fluctuations (red) in the inertial (power law) range. Azimuthal velocity fluctuations (green) have the least amplitude throughout the wavenumber range.
@@ 4.7 Reynolds stresses
was found to be almost constant at $\approx 0.2$-$0.3$ throughout the domain. This suggests that the shapes of the structures responsible for $T_{r\phi}$ are similar at all locations of the domain; only their intensity and/or size varies. The trace $T_{kk}$ (not shown), which is twice the kinetic energy, peaks at the midplane but maintains high values up to $z\approx 1.5H_{0}$. Strong peaks in all the stresses near the left and right radial boundaries are considered to be artifacts due to boundary conditions.
@@ 4.8 The turbulence $\alpha$ parameter
An explanation one might initially suggest for the increase is the fact that the appropriate velocity scale for $\nu_{\mathrm{t}}$ should be the driving vertical shear $\Delta u_{\phi}(r)=|u_{\phi}(r,z=H)-u_{\phi}(r,z=0)|$ rather than the sound speed as assumed by Shakura & Sunyaev (1973). Unfortunately, however, $\Delta u_{\phi}(r)\propto c_{\mathrm{i}}(r)$ for the VSI basic state. Another explanation one might suggest is that the radial wavelength $\ell_{r}(r)$ increases with $r$ faster than $H(r)$ does (§4.4).
Figure 19 shows that with increasing resolution (starting with the brown curve) $\alpha$ initially rapidly increases to the green curve but then decreases to the blue and black curves. Inspection of the vorticity field for the brown case reveals minimal radial waviness of shear layers, which would contribute to a suppression of $T_{\phi r}$. Why should the highest resolution case (black curve) have lower $\alpha(r)$ than the less resolved blue and green cases? We conjecture that large scale waviness present at lower resolution breaks down into smaller vortices at higher resolution, leading to a smaller $T_{r\phi}$ correlation.
@@ 5 Late time artifacts in the axisymmetric simulation
Given that flow length scales increase with time in our simulation (see below) as well as in the simulations of M24a, i.e., that there is an inverse cascade, the time to reach stationarity in any given simulation would depend on the radial domain size.
@@
Figure 23 shows the perturbation azimuthal and vertical vorticity components at $t/T_{0}=197.6$, which is at a late time but prior to the stationary state. Their structure is very different from the 3D case and what is observed at early times in the axisymmetric case. The dominant component now is $\delta\omega_{z}$ rather than $\omega_{\phi}$, which translates to a larger magnitude of the azimuthal velocity perturbation compared to the vertical velocity. The shear layers of $\omega_{\phi}$ that formed the edges of vertical jets at earlier times are still present as is the pattern $\omega_{\phi}\lessgtr 0$ for $z\gtrless 0$ within these shear layers. However, the flow between the shear layers consists of large eddies and is reminiscent of 2D turbulence. Panel (b) shows that the $\approx 7$ layers of vertical vorticity seen at $t/T_{0}=36.1$ (Figure 5) have now intensified, and only 4 remain. At a later time ($t/T_{0}=403.7$, not shown) only 2 such layers remain. In other words, flow length scales increase with time. The layers of strong $\delta\omega_{z}$ are associated with jumps in the total specific angular momentum $j_{z}=u_{\phi}r$ (panel (c)), which has a strong vertically coherent staircase profile with respect to $r$.
@@
This can be seen more clearly in Figure 24, which shows the midplane profiles of $j_{z}(r)$ at $t/T_{0}=197.6$ and $t/T_{0}=403.7$. The latter is in the stationary state and has only 2 jumps in $j_{z}$. As shown earlier (Figure 4, blue line), in 3D there is no large scale rearrangement of angular momentum, rather only a weak one in the upper layers of the disk. The circulation, $\Gamma_{z}=2\pi j_{z}$, of circular material lines is proportional to $j_{z}$. Hence by Stokes’ theorem, a region of nearly constant $\Gamma_{z}(r)$ must be irrotational. For inviscid barotropic flow, Kelvin’s theorem asserts that that $D\Gamma_{z}/Dt=0$ for each circular material line. Hence there are three possible mechanisms for the formation of the staircase structure: vertical transport of $j_{z}(z)$ given its vertical gradient (this being the mechanism that operates at early times), radial mixing of $j_{z}(r)$ by eddies between the shear layers, or baroclinic torque. For a circular material line we have that
@@
where $A$ is the circular surface enclosed by the circle and $\widehat{\mathbf{z}}$ is the unit vector normal to the surface. However, the $z$-component of the baroclinic term in the integrand is zero for axisymmetric flow. The vertical gradient of $j_{z}$ is weaker than its radial gradient, and radial velocity fluctuations are larger than vertical ones at late times. Hence, we conclude that radial mixing of $j_{z}$ by eddies in regions bounded by vertical shear layers is responsible for the staircase profile of $j_{z}$ at late times.
@@
In the axisymmetric simulations of MF24a there is also a clear difference between early and late times. Let us focus on their dg3c4_2048 simulation. Their Figure 1 (left hand plots) and Figure 2 (top) shows very many small steps in the $j_{z}$ staircase at an earlier time. On the other hand, their Figure 11 shows the same run at 1360 orbits. Clearly, an increase of length scales has taken place leading to only 3 jumps in $j_{z}$.
@@ 6.1 Caveats for applying RWI to VSI simulations
following fluid particles. Here $P$ is the vertically integrated pressure, $\Sigma$ is the surface density, and $\gamma$ is the adiabatic exponent. A necessary condition for RWI is that, in the basic state, the function
@@
has a local maximum. The thin disk assumption is valid only for motions whose horizontal length scale $\ell\gg H(r)$. On the other hand, the radial size of LPVs observed in simulations is about $H(r)$. In addition, the locally isothermal equation of state, $p=\rho c_{\mathrm{i}}^{2}(r)$, assumed in the present and other VSI simulations implies that radial motions fail to conserve $S$ following fluid particles. Likewise, it remains to be shown whether radiative hydrodynamic VSI simulations that invoke RWI (e.g., Flock et al., 2020) possess an effective adiabatic exponent. Despite the lack of adiabaticity, previous works plot a vertically averaged $\mathcal{L}(r)$, presumably setting $\gamma=1.4$. In spite of these caveats, it appears that the presence of LPVs in VSI simulations is correlated with some (but not all) local maxima in $\mathcal{L}(r)$ and we will discuss these simulations in the next subsection.
@@ 6.4 Speculation
Why do M20 observe LPVs in the interior of the domain for the same disk parameters but we do not? The answer may lie in the link, discussed by MF24b, between maxima in $\mathcal{L}(r)$ and the $j_{z}(r)$ staircase that is prominent in axisymmetric $j_{z}(r)$. Specifically, a region of flattened $j_{z}(r)=u_{\phi}r$ corresponds to more rrotational flow ($u_{\phi}\approx\mathrm{const}/r$) and therefore suppressed $\omega_{z}$. Since mass also accumulates in these regions, a local maximum is created in the first factor ($\Sigma/2\omega_{z}$) of $\mathcal{L}$. Therefore, one possibility is that M20 has a stronger $j_{z}(r)$ staircase in 3D than the present simulation, which leads to several $\mathcal{L}(r)$ maxima in the interior of the domain. It is tempting to attribute this to M20’s lower resolution; however, our low resolution simulations also failed to produce LPVs away from boundaries. Further effort is required to ascertain the reasons for this difference between our and the M20 simulation.
@@ 7 Closing remarks
    Radial wavelength. In axisymmetric VSI, the radial wavelength at early times corresponds to the most amplified linear mode, but then inexorably increases in time. In 3D, the ratio
    $\frac{\ell_{r}(r)}{\lambda_{r,\mathrm{max}}(r)}=\left[\pi|q|(H_{0}/R_{0})\right]^{-1}\frac{\ell_{r}(r)}{H(r)}$ (61)
    of the radial wavelength, $\ell_{r}(r)$, to the most linearly amplified wavelength ($\lambda_{r,\mathrm{max}}$) increases from 2.4 to 4.4 as $r$ increases from $8.6H_{0}$ to $11.6H_{0}$.
@@
    Layers of positive $\delta\omega_{z}$. Axisymmetric VSI has coherent vertical layers of $\delta\omega_{z}>0$ that increase in strength with $|z|$ and are coincident with interfaces between the upward and downward jets (Figure 5). Their formation and the preference for $\delta\omega_{z}>0$ can be explained by the vertical transport of specific angular momentum $j_{z}\equiv u_{\phi}r$ by the vertical jets. This phenomenon is equivalent to the formation of a staircase profile of $j_{z}(r)$ first discovered by Klahr et al. (2023) and MF24a,b in their axisymmetric simulations. We find that these layers are also present in 3D but with much weaker jumps in specific angular momentum compared to the axisymmetric case (Figure 4(b)). These layers become stronger and more azimuthally coherent with increasing $|z|$ and undergo KH-like instability (Figure 13a), likely modified by background shear and rotation. The resulting vortices are small and possess all three vorticity components.
@@
    Variation of $\alpha$ with $r$. There is a linear increase of $\alpha(r)$ with radius with a slope (at the highest resolution) of about 14% the midradius value per scale height. This increase currently lacks an explanation; it may be related to the increase of radial spacing, $\ell_{r}(r)$, with $r$ that is faster than $H(r)$ (see item 1, which itself requires an explanation).
@@
    Late time artifacts in axisymmetric VSI. As discussed above, some features of the 3D simulation can be understood by appealing to physics in the early stage of the axisymmetric simulation. However, at late times, the axisymmetric simulations have artifacts that are not present in 3D. These include increasing flow length scales and a strong staircase pattern in the specific angular momentum $j_{\phi}=u_{\phi}r$ with an accompanying redistribution of mass.
We are grateful to Dr. J. David Melon Fuksman (Max Planck Institut für Astronomie, Heidelberg) for his detailed comments, to Drs. Paul Estrada and Ted Manning at NASA Ames for performing a careful internal review, and to the referee for many useful suggestions that led to an improved manuscript. OMU acknowledges partial support from NASA’s TCAN (Theoretical and Computational Astrophysical Networks) grant #80NSSC21K0497.